### v3

@@ Abstract
Control-flow attestation unifies the worlds of control-flow integrity and platform attestation by measuring and reporting a target’s run-time behaviour to a verifier. Trust assurances in the target are provided by testing whether its execution follows an authorised control-flow path. The problem has been explored in various settings, such as assessing the trustworthiness of cloud platforms, cyber-physical systems, and Internet of Things devices. Despite a significant number of proposals being made in recent years, the area remains fragmented, addressing different adversarial behaviours, verification paradigms, and deployment challenges. In this paper, we present the first survey of control-flow attestation, examining the core ideas and solutions in state-of-the-art schemes. In total, we survey over 30 papers published between 2016–2024, consolidate and compare their key features, and pose several challenges and recommendations for future research in the area.
@@ 1 Introduction
In recent years, a large number of CFA schemes have been proposed in the literature, focussing on different measurement approaches, roots of trust, adversarial models, and hardware and software assumptions. For instance, different hardware security technologies have been suggested for protecting $\mathcal{P}$’s measurement and reporting procedures [4, 5, 6, 7]; various static and dynamic methods have been proposed for program instrumentation [4, 8, 9, 10, 11]; and numerous control-flow transfer types have been monitored [12, 7, 13, 10, 14]. Despite receiving significant attention from the research community, no survey exists that consolidates and reflects on the state of the art. In this paper, we address this gap through a comprehensive review of existing CFA proposals and make the following contributions:
  • We present the first systemisation of knowledge paper on control-flow attestation, covering over 30 CFA proposals over approximately the last decade since CFA was first introduced in the literature.
@@ 1.1 Related Work
To the best of our knowledge, no extensive survey currently exists summarizing CFA schemes; however, several existing papers survey remote attestation (RA) more generally in different contexts, which we summarise in Table 1.
In 2016, Steiner & Lupu [15] presented one of the first surveys covering traditional remote attestation approaches applicable to wireless sensor networks. Arias et al. [16] (2018) provided a short review of device attestation methods for IoT devices, touching principally on software, hardware, and swarm-based attestation in the static measurement model. In 2020, Sfyrakis & Groß [17] reviewed different hardware approaches for attesting devices in network infrastructures, such as those used in infrastructure-as-a-service solutions. In 2021, Johnson et al. [18] proposed a taxonomy of RA schemes for embedded systems, categorising proposals according to their root of trust, evidence type, evidence gathering, packaging and verification, and scalability. Ankergård et al. [19] examined software-only remote attestation schemes without relying on hardware-based roots of trust for IoT devices. Kuang et al. [20] (2022) survey remote attestation schemes for IoT devices in general, detailing attestation proposals targeting different attacks (e.g. static and run-time software attacks, and physical attacks). Lastly, Ambrosin et al. [21] review solutions in the area of collective remote attestation, where multiple devices are attested as a whole rather than individually.
@@
Work Year Description
Steiner & Lupu [15] 2016 RA methods for wireless sensor networks.
Arias et al. [16] 2018 A brief review of device attestation methods.
Sfyrakis & Groß [17] 2020 Hardware methods for attesting network infrastructures.
Johnson et al. [18] 2021 A review of RA schemes for embedded systems.
Ankergård et al. [19] 2021 State-of-the-art software-based RA for IoT devices.
Kuang et al. [20] 2022 A general review of RA schemes for IoT devices.
Ambrosin et al. [21] 2022 Collective RA schemes for large IoT networks.
This Paper 2024 A comprehensive analysis of control-flow attestation.
In some of these surveys [16, 18, 21], control-flow attestation is described briefly but not comprehensively, considering only early control-flow attestation schemes (e.g. [4, 8]). They do not present a dedicated treatment to the large number of CFA proposals that have emerged over the past decade. We directly address this gap in this paper. Survey papers on control-flow integrity more generally have also been published on verification methods [22] and hardware-based CFI systems [23]. This paper offers a comprehensive treatment of proposals that attest devices using run-time control-flow behaviours. Specifically, we consider proposals that contain mechanisms for securely measuring and reporting the state of a proving device. Proposals that focus on control-flow integrity, but do not provide attestation, are considered useful supplementary literature; however, they are orthogonal to this work.
@@ 2.1.2 Non-Control Data Attacks
In Figure 1, we give some common points in which control- and data-flow attacks may be used. In the literature, adversaries with the ability to perform data-oriented attacks are considered to be highly sophisticated [32, 34, 35]. We emphasise that, if no additional context-sensitivity is introduced, or additional checks are performed on critical data variables, then it is possible for non-control-data and DOP attacks to bypass control-flow integrity mechanisms.
@@ 2.2 Control-Flow Transfer Types
In CFI and CFA, monitoring different control-flow transfer types is used to detect particular unauthorised paths taken by a program. It is often undesirable to monitor every possible control-flow transfer due to the significant performance overhead that this can occur, especially on constrained devices. Indeed, not all control-flow transfers are relevant to the security of a system, and it is intuitive to monitor transfers that are common targets for particular exploits of interest. For example, one may wish to pay particular attention to indirect jumps and calls in order to address ROP and JOP attacks. In Figure 2, we give some examples of such transfers used by X86-64 platforms. We now describe the control-flow transfer types which have been used in existing work:
{minted}
[frame=lines,framesep=2mm]nasm ;; Conditional direct jump (jump if equal) cmp rax, rbx ; Compare RAX and RBX je mylabel ; Jump to ’mylabel’ if RAX = RBX ;; Unconditional direct jumps jmp mylabel ; Jump to ’mylabel’ jmp 0x00401000 ; Jump to the address 0x00401000 ;; Direct calls call myfunc ; Call the function ’myfunc’ call 0x00401234 ; Call the function at 0x00401234 ;; Indirect jumps jmp rax ; Jump to an address in register RAX jmp [0x00401234] ; Jump to an address pointer 0x00401234 jmp [rbx+rcx*4] ; Register-indirect jump with an offset jmp [rsp] ; Jump to an address on the stack ;; Indirect calls call rax ; Call an address in RAX call [0x00401234] ; Call a function pointer at 0x00401234 call [rbx+4] ; Register-indirect call with an offset call [rsp] ; Use the stack to call a function ;; Return instructions ret
@@
    All control-flow transfers: This is the most expansive definition. We assume that $\mathcal{C}$ is a maximal set, or has the greatest cardinality, for the target architecture. In practice, all possible control-flow instructions are examined—conditional jumps, direct calls, indirect calls, etc.—imposing the greatest overhead during monitoring and measurement. It is also the most invasive, requiring the greatest level of instrumentation of the target program (see §6). Monitoring all control-flow transfers has been used in many current proposals [37, 12, 34, 38, 36, 39, 40, 9, 41].
@@
Some proposals apply different CFI policies depending on the location of the code being examined with the aim of reducing performance overhead. For example, by dividing code into critical and non-critical functions through manual policy specification or using statistical and machine learning-based models [11, 9, 5, 42]. For example, ARI [43] divides code into critical and non-critical compartments based on a policy set. All control-flow transfers are monitored except unconditional direct jumps within critical compartments, but transfers within non-critical areas are not considered. OAT [6] attests partial code regions that are specified (annotated) by programmers. Within the attested code, conditional direct jump, indirect jump, indirect call, and return instructions are covered. ACFA [44] monitors all control-flow transfers within pre-determined code regions including any transfer from outside to inside the region. Other approaches monitor all control-flow transfers within areas that process sensitive data, rather than denoted code regions (see DIAT [45]). Tracking load/store operations on critical data in memory has also been used, thus bridging the areas of control flow and data flow, but this remains a nascent area of research [34, 45].
@@ 3 Introducing Control-Flow Attestation
Reporting] [[Hybrid
Attestation]] [Verifiable
@@
Execution] [Hybrid] [Path Log]] [Hardware] ] [Adversarial
Behaviour [Non-control
Data Attacks [DOP]] [Code
Injection [Static] [Run-time]] [Code
Reuse [COP] [JOP] [ROP]] ] ]
@@ 4.2 Continuous Reporting
4.3 Hybrid Attestation
We define hybrid attestation as a combination of local monitoring and remote attestation. Geden and Rasmussen [47] follow the same offline stage as conventional remote attestation where $\mathcal{V}$ extracts a CFG, known as a runtime integrity model, before deploying the proving device. The model is stored internally in a discrete hardware security module (HSM) aboard $\mathcal{P}$. The HSM continuously takes control-flow events as input from the CPU in real-time in order to thwart time-of-check-to-time-of-use (TOCTOU) problem in conventional remote attestation. A remote verifier can acquire a report from the HSM, showing whether any security policy violations were detected over the course of the target program’s execution.
@@ 4.5 Collective Attestation
Figure 8: System overview of DIAT [45]. Unlike conventional schemes, measurements from each sub-system’s CFGs are aggregated within attestation reports.
@@ 5 Trust Anchors and Assumptions for Control-Flow Attestation
The trust anchor is commonly used to protect $\mathcal{P}$’s measurement procedure and any key material used in generating attestation reports. For some proposals, this requires that the trust anchor passes traditional static attestation and that it has dedicated memory to store material that only the anchor itself can access [41, 37, 13, 34]. For proposals that rely on TEEs, it is assumed that potentially applicable high-cost attacks cannot be performed by an adversary, such as hardware fault injections and side-channel attacks [51]. TEE-based proposals sometimes place trust in certain operating system components, such as kernel modules that initiate the communication with the TEE [38, 6]. In proposals where the kernel serves as $\mathcal{P}$’s trust anchor, it is somtimes assumed that a method for protecting cryptography keys is available, such as a TPM or a proprietary facility (see Intel MPK in [58]).
Very few CFA proposals are purely based in software, and most systems have some degree of hardware dependency. We observe that CFA schemes can be split into three categories based principally on where security-critical CFA components are implemented on $\mathcal{P}$:
@@
  • Hardware-based proposals rely on a dedicated hardware for performing the measurement and reproting operations. This may be a discrete on-chip security module, an off-chip micro-controller, or a set of bespoke CPU extensions.
  • Hardware-assisted systems use software whose security is underpinned by a hardware root of trust. TEE-based proposals are characterised as such, as well as proposals that rely on read-only memory (ROM), secure boot [45], and a TPM [13] in tandem with a normal operating system.
@@ 5.1 Trusted Execution Environments
    The trampoline forwards the control-flow event to the TEE over a secure monitor that validates the data and performs the world switch to the TEE.
@@ 5.2 Normal World Schemes
A small number of proposals deploy the process components, $\mathcal{M}$ and $\mathcal{R}$, in the same execution environment as the target program, whether in kernel- or user-space. This may be considered the ‘normal’ or ‘non-secure’ world, such as Android, Linux or a real-time operating system (RTOS). These schemes do not require a trampoline component to forward control-flow instances to the measurement module at runtime. ScaRR [13] relies on kernel-space modules for measuring relevant control-flow events and creating attestation reports. User-space programs hook relevant control-flow transfers, which are passed to the kernel modules through a user-space library (Figure 10). The design decision was motivated by performance. It is stated that using a TEE (Intel SGX) imposed considerable overhead “due to SGX clearing the Translation-Lookaside Buffer (TLB)…at each enclave exit. This caused frequent page walks after each enclave call.” (p.8, [13]). Tiny-CFA [62], a hardware-based CFA scheme designed for low-end microcontroller units (MCUs), makes a similar argument.
@@ 5.3.1 Custom Extensions
CHASE [35] provides a configurable hardware design capable of detecting different adversaries. A selection of up to 10 sub-modules can be used to detect various control- and data-flow attacks, such as DOP and DDM attacks. To achieve this, the modules hook into the fetch and decode phases of the processor pipeline, which outputs features to a hashing module that continuously generates hash measurements for attestation reports. Depending on the configuration, both local and remote attestation can be provided. Addressing the strongest adversarial models, e.g. DOP attacks, requires the use of all 10 sub-modules, whereas the simplest threats requires the use of only four sub-modules.

### v4

@@ Abstract
Control-flow attestation unifies the worlds of control-flow integrity and platform attestation by measuring and reporting a target’s run-time behaviour to a verifier. Trust assurances in the target are provided by testing whether its execution follows an authorised control-flow path. The problem has been explored in various settings, such as assessing the trustworthiness of cloud platforms, cyber-physical systems, and Internet of Things devices. Despite a significant number of proposals being made in recent years, the area remains fragmented, with different adversarial behaviours, verification paradigms, and deployment challenges being addressed. In this paper, we present the first survey of control-flow attestation, examining the core ideas and solutions in state-of-the-art schemes. In total, we survey over 30 papers published between 2016–2024, consolidate and compare their key features, and pose several challenges and recommendations for future research in the area.
@@ 1 Introduction
In recent years, a large number of CFA schemes have been proposed in the literature, focusing on different measurement approaches, roots of trust, adversarial models, and hardware and software assumptions. For instance, different hardware security technologies have been suggested for protecting $\mathcal{P}$’s measurement and reporting procedures [4, 5, 6, 7]; various static and dynamic methods have been proposed for program instrumentation [4, 8, 9, 10, 11]; and numerous control-flow transfer types have been monitored [12, 7, 13, 10, 14]. Despite receiving significant attention from the research community, no survey exists that consolidates and reflects on the state of the art. In this paper, we address this gap through a comprehensive review of existing CFA proposals and make the following contributions:
  • We present the first systemisation of knowledge paper on control-flow attestation, covering 31 CFA proposals over approximately the last decade since CFA was first introduced in the literature.
@@ 1.1 Related Work
To the best of our knowledge, no extensive survey currently exists summarising CFA schemes; however, several existing papers survey remote attestation (RA) more generally in different contexts, which we summarise in Table 1.
In 2016, Steiner & Lupu [15] presented one of the first surveys covering traditional remote attestation approaches applicable to wireless sensor networks. Arias et al. [16] (2018) provided a short review of device attestation methods for IoT devices, touching principally on software, hardware, and swarm-based attestation in the static measurement model. In 2020, Sfyrakis & Groß [17] reviewed different hardware approaches for attesting devices in network infrastructures, such as those used in infrastructure-as-a-service solutions. In 2021, Johnson et al. [18] proposed a taxonomy of RA schemes for embedded systems, categorising proposals according to their root of trust, evidence type, evidence gathering, packaging and verification, and scalability. Ankergård et al. [19] examined software-only remote attestation schemes without relying on hardware-based roots of trust for IoT devices. Kuang et al. [20] (2022) survey remote attestation schemes for IoT devices in general, detailing attestation proposals targeting different attacks (e.g. static and run-time software attacks, and physical attacks). Ambrosin et al. [21] review solutions in the area of collective remote attestation, where multiple devices are attested as a whole rather than individually. Recently, Ammar et al. [22] (2024) provided a summary of runtime integrity mechanisms, focusing on CFI and a selection of CFA papers, discussing their commonalities and distinctions, such as architecture, scalability, and attack vectors. However, their work does not offer a comprehensive review of state-of-the-art CFA schemes, which is the focus of our analysis.
@@
Work Year Description # CFA Papers
Steiner & Lupu [15] 2016 RA methods for wireless sensor networks. 1
Arias et al. [16] 2018 A brief review of device attestation methods. 2
Sfyrakis & Groß [17] 2020 Hardware methods for attesting network infrastructures. 0
Johnson et al. [18] 2021 A review of RA schemes for embedded systems. 6
Ankergård et al. [19] 2021 State-of-the-art software-based RA for IoT devices. 2
Kuang et al. [20] 2022 A general review of RA schemes for IoT devices. 8
Ambrosin et al. [21] 2022 Collective RA schemes for large IoT networks. 5
Ammar et al. [22] 2024 A review of runtime integrity mechanisms 18
This Paper 2024 A comprehensive analysis of control-flow attestation. 31
In several surveys [16, 18, 21, 22], control-flow attestation is described briefly, focusing on early control-flow attestation schemes (e.g. [4, 8]). Survey papers on control-flow integrity more generally have also been published on verification methods [23] and hardware-based CFI systems [24]. To the best of our knowledge, we are not aware of a dedicated treatment of CFA proposals that have emerged in recent years. We directly address this gap in this paper.
1.2 Scope
To scope this work, we focus on the problem of attesting a target platform by collecting evidence derived from control-flow events. We introduce definitions based on the principles and definitions established in [25].
Definition 1.1 (Verifier and Target).
A verifier is an entity, such as a computer on a network, that makes a decision about another party. A target or prover is a party, such as an untrusted device, about which a verifier must make such a decision.
The verifier checks whether the target system state meets certain security requirements by analyzing evidence provided by the target. This evidence serves as proof in a process known as attestation, which helps the verifier determine if the target is free from unauthorised changes, malware, or other violations that may compromise security.
Definition 1.2 (Attestation and Verification).
Attestation is the activity of making a claim to a verifier about the properties of a target by supplying evidence which supports that claim. Verification refers to the verifier’s decision-making process based on the information it receives during the attestation process from the prover.
Definition 1.3 (Measurement).
A proving device collects evidence about a target to be attested through direct and local observation using its measurement process. The verifier uses its verification process to evaluate this evidence.
Substantial work has been conducted in generating evidence using static measurements, such as cryptographic hashes of firmware or software binaries [17, 18, 20, 15, 19, 16, 21]. While static attestation provides a snapshot of the target’s state at a given point in time, it lacks the ability to detect run-time attacks, such as control-flow hijacking (see §2.1), which may occur after the target is under execution and been deemed trustworthy.
To address these limitations, the area of control-flow attestation has emerged, where evidence is collected at run-time during the target’s execution from its control-flow behaviour. In this work, we conducted an extensive literature search using multiple databases, such as IEEE Xplore, ACM Digital Library, and Google Scholar. We used keywords and phrases closely associated with the topic area, including “run-time”, “control-flow”, “dynamic” and “hybrid” with the terms “attestation”, “trust”, and “trusted computing”. We also incorporated forward and backward citation tracking to identify additional relevant works, particularly using related survey papers in Table 1. Papers were selected if they satisfied the following criteria:
  • The work proposed a new method of attestation aimed at establishing trust in a target entity, involving a prover and a verifier.
  • The proposal incorporated control-flow information produced at run-time by the target (prover) within an attestation response during the target’s execution.
As such, we are agnostic to the particular instruction set architecture, manufacturer or device under consideration. Proposals that focus on control-flow integrity, but not attestation, are considered useful supplementary literature (see [26]). However, such proposals are ultimately considered outside the scope of this work. Similarly, static attestation methods, and those that rely on a combination of static attestation and run-time integrity enforcement, e.g. [27], are also excluded.
@@ 2.1.2 Non-Control Data Attacks
In Figure 1, we give some common points in which control- and data-flow attacks may be used. In the literature, adversaries with the ability to perform data-oriented attacks are considered to be highly sophisticated [36, 38]. We emphasise that, if no additional context-sensitivity is introduced, or additional checks are performed on critical data variables, then it is possible for non-control-data and DOP attacks to bypass control-flow integrity mechanisms.
@@ 2.2 Control-Flow Transfer Types
In CFI and CFA, monitoring different control-flow transfer types is used to detect particular unauthorised paths taken by a program. It is often undesirable to monitor every possible control-flow transfer due to the significant performance overhead that this can occur, especially on constrained devices. Indeed, not all control-flow transfers are relevant to the security of a system, and it is intuitive to monitor transfers that are common targets for particular exploits of interest. For example, one may wish to pay particular attention to indirect jumps, indirect calls, and return in order to address ROP and JOP attacks. In Figure 2, we give some examples of such transfers used by X86-64 platforms. [^4] X86-64 is chosen for illustration purposes; similar examples apply on other instruction set architectures, including ARM and RISC-V. We observe that measurements in CFA proposals are modelled as BBLs or control-flow transfers, rather than specific assembly instructions. We now describe the control-flow transfer types which have been used in existing work:
@@
    All control-flow transfers: This is the most expansive definition. We assume that $\mathcal{C}$ is a maximal set, or has the greatest cardinality, for the target architecture. In practice, all possible control-flow instructions are examined—conditional jumps, direct calls, indirect calls, etc.—imposing the greatest overhead during monitoring and measurement. It is also the most invasive, requiring the greatest level of instrumentation of the target program (see §6). Monitoring all control-flow transfers has been used in many current proposals [40, 12, 38, 41, 39, 42, 43, 44].
@@
Some proposals apply different CFI policies depending on the location of the code being examined with the aim of reducing performance overhead. For example, by dividing code into critical and non-critical functions through manual policy specification or using statistical and machine learning-based models [11, 9, 5, 45]. For example, ARI [46] divides code into critical and non-critical compartments based on a policy set. All control-flow transfers are monitored except unconditional direct jumps within critical compartments, but transfers within non-critical areas are not considered. OAT [6] attests partial code regions that are specified (annotated) by programmers. Within the attested code, conditional direct jump, indirect jump, indirect call, and return instructions are covered. ACFA [47] monitors all control-flow transfers within pre-determined code regions including any transfer from outside to inside the region. Other approaches monitor all control-flow transfers within areas that process sensitive data, rather than denoted code regions (see DIAT [48]). Tracking load/store operations on critical data in memory has also been used, thus bridging the areas of control flow and data flow, but this remains a nascent area of research [38].
@@ 3 Introducing Control-Flow Attestation
Reporting] [[Local Verification
with Remote Reporting]] [Verifiable
@@
Execution] [Hybrid] [Path
Log]] [Whitelist-
based] ] [Adversarial
Behaviour [Control Data
Attacks [Code
Injection] [Code
Reuse [COP] [JOP] [ROP] ] ] [Non-control
Data Attacks [DDM [Branch
Variables] [Loop
Variables] ] [Data-oriented
Attacks [DOP] ] ] ] ]
@@ 4.2 Continuous Reporting
4.3 Local Verification with Remote Reporting
A separate class of CFA proposals we have identified is those that use a combination of local verification and remote reporting. Geden and Rasmussen [50] follows the same offline stage as conventional remote attestation where $\mathcal{V}$ extracts a CFG, known as a runtime integrity model, before deploying the proving device. The model is stored internally in a discrete hardware security module (HSM) aboard $\mathcal{P}$. The HSM continuously takes control-flow events as input from the CPU in real-time and performs CFI check locally. A remote verifier can acquire a report from the HSM, showing whether any security policy violations were detected over the course of the target program’s execution.
@@ 4.5 Collective Attestation
Figure 8: System overview of DIAT [48]. Unlike conventional schemes, measurements from each module’s CFG are aggregated within attestation reports.
@@ 5 Trust Anchors and Assumptions for Control-Flow Attestation
The trust anchor is commonly used to protect $\mathcal{P}$’s measurement procedure and any key material used in generating attestation reports. For some proposals, this requires that the trust anchor passes traditional static attestation and that it has dedicated memory to store material that only the anchor itself can access [44, 40, 13, 38]. For proposals that rely on TEEs, it is assumed that potentially applicable high-cost attacks cannot be performed by an adversary, such as hardware fault injections and side-channel attacks [55]. TEE-based proposals sometimes place trust in certain operating system components, such as kernel modules that initiate the communication with the TEE [41, 6]. In proposals where the kernel serves as $\mathcal{P}$’s trust anchor, it is sometimes assumed that a method for protecting cryptography keys is available, such as a TPM or a proprietary facility (see Intel MPK in [62]).
Very few CFA proposals are purely based in software, and most systems have some degree of hardware dependency. We observe that CFA schemes can be split into two categories based principally on where security-critical CFA components are implemented on $\mathcal{P}$:
@@
  • Hardware-based proposals rely on a dedicated hardware for performing the measurement and reporting operations. This may be a discrete on-chip security module, an off-chip micro-controller, or a set of bespoke CPU extensions. Alternatively, the proposal uses software whose security is underpinned by a hardware root of trust, such as TEE-based proposals and those that rely on read-only memory (ROM), secure boot [48], or a TPM [13].
@@ 5.1 Trusted Execution Environments
    The trampoline forwards the control-flow event to the TEE over a secure monitor that performs the world switch to the TEE.
@@ 5.2 Normal World Schemes
A small number of proposals deploy the process components, $\mathcal{M}$ and $\mathcal{R}$, in the same execution environment as the target program, whether in kernel- or user-space. This may be considered the ‘normal’ or ‘non-secure’ world, such as Android, Linux or a real-time operating system (RTOS). These schemes do not require a trampoline component to forward control-flow events to the measurement module at runtime. ScaRR [13] relies on kernel-space modules for measuring relevant control-flow events and creating attestation reports. User-space programs hook relevant control-flow transfers, which are passed to the kernel modules through a user-space library (Figure 10). The design decision was motivated by performance. It is stated that using a TEE (Intel SGX) imposed considerable overhead “due to SGX clearing the Translation-Lookaside Buffer (TLB)…at each enclave exit. This caused frequent page walks after each enclave call.” (p.8, [13]). Tiny-CFA [66], a hardware-based CFA scheme designed for low-end microcontroller units (MCUs), makes a similar argument.
@@ 5.3.1 Custom Extensions
CHASE [53] provides a configurable hardware design capable of detecting different adversaries. A selection of up to 10 sub-modules can be used to detect various control-flow attacks, such as code-reuse attacks and DDM attacks. To achieve this, the modules hook into the fetch and decode phases of the processor pipeline, which outputs features to a hashing module that continuously generates hash measurements for attestation reports. Depending on the configuration, both local and remote attestation can be provided. Addressing the strongest adversarial models, e.g. control-data and non-control-data attacks, requires the use of 9 sub-modules, whereas the simplest threats requires the use of only four sub-modules.